NOAA SATOPS Morning Report: Wednesday, January 23, 2013


SATOPS Morning Report: Wednesday, January 23, 2013

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 023

GOES
GOES-12 intermittent image frame breaks due to BER at WCDA: 12:08z to 12:23z, 20:40z to 20:49z, 01:47Z, AND 05:54Z.

GOES-15 image frame breaks from 01:39z to 02:00z

All other GOES operations were nominal over the past 96 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 96 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 96 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 96 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
MET 10: Missing and partial Meteosat-10 FD frames from 1300z to 1345z.

GINI West AWIPS lost 4 sectors at 0130z and the 0140z was also lost due to high BER at Wallops. The 0200z was skipped due to delays in processing the 0145z data. The GINI East AWIPS lost 2 sectors of the 0200z.

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.